# Data-Driven Spectral Analysis Through Pseudo-Resolvent Koopman Operator in Dynamical Systems

### v1

@@ Abstract
We present a data-driven method for spectral analysis of the Koopman operator based on direct construction of the pseudo-resolvent from time-series data. Finite-dimensional approximation of the Koopman operator, such as those obtained from Extended Dynamic Mode Decomposition, are known to suffer from spectral pollution. To address this issue, we construct the pseudo-resolvent operator using the Sherman-Morrison-Woodbury identity whose norm serves as a spectral indicator, and pseudoeigenfunctions are extracted as directions of maximal amplification. We establish convergence of the approximate spectrum to the true spectrum in the Hausdorff metric for isolated eigenvalues, with preservation of algebraic multiplicities, and derive error bounds for eigenvalue approximation. Numerical experiments on pendulum, Lorenz, and coupled oscillator systems demonstrate that the method effectively suppresses spectral pollution and resolves closely spaced spectral components.
@@ 1 Introduction
The study of nonlinear dynamical systems is central to many scientific and engineering problems [ansov1969geodesic, birkhoff1927dynamical, katok1995introduction, poincare1890probleme, smale1967differentiable, strogatz2024nonlinear]. Traditionally, their analysis relies on first-principles models together with tools such as phase-space geometry [poincare1890probleme], perturbation methods [kato1995perturbation] for weak nonlinearities, and direct numerical simulation. These approaches have delivered deep insight, but they can be difficult to use when an accurate governing model is unavailable, when model parameters are uncertain, or when the goal is to infer global dynamical features from observations rather than describe local behavior near equilibria [hartman1960lemma, hartman2002ordinary]. Meanwhile, modern experiments and simulations increasingly produce large amount of time-series measurements, which has attracted growing interest in data-driven methods that aim to characterize dynamical behavior directly from data.
@@
A natural perspective for data-driven analysis is to focus on the evolution of observable quantities rather than the underlying state variables. In many applications, one does not have direct access to the full state of a system but instead observations of the state, e.g., temperatures, velocities, concentrations, or other quantities of interest. The Koopman operator provides a mathematical framework for this viewpoint; more specifically, it describes how observables evolve in time by composing them with the flow map of the dynamical system [Koopman1931, koopman1932dynamical, mezic2005spectral]. Although the underlying dynamics may be nonlinear, the Koopman operator acts linearly on the space of observables, enabling spectral analysis to reveal dynamical features such as oscillation frequencies, growth and decay rates, and coherent structures. Dynamic Mode Decomposition (DMD) [tu2013dynamic] and its extension, Extended Dynamic Mode Decomposition (EDMD) [Williams_2015], provide practical algorithms for approximating the Koopman operator and its spectral properties directly from time-series data, and theoretical foundations have established convergence of these approximations under appropriate conditions [colbrook2024rigorous, Ishikawa2025KoopmanIntrinsicRiggedRKHS, KLUS2020132416, korda2018convergence, xu2025datadrivenframeworkkoopmansemigroup].
@@
Despite these convergence guarantees, a fundamental difficulty arises when extracting spectral information from finite-dimensional approximations: the phenomenon of spectral pollution [colbrook2024rigorous]. The approximating matrix is finite-dimensional and therefore possesses a discrete spectrum, whereas the true Koopman operator may have continuous spectral components including continuous spectrum [lusch2018deep]. Consequently, computed eigenvalues may include spurious points that do not correspond to any part of the true spectrum, in other words, these are drawbacks of the finite-dimensional truncation rather than intrinsic dynamical features. Conversely, true spectral points may be missed entirely if they are not well-represented by the chosen dictionary. This problem is not merely a matter of insufficient data or dictionary size; it is a structural consequence of approximating an infinite-dimensional operator by finite-dimensional matrices .
@@
To address spectral pollution, Residual Dynamic Mode Decomposition [colbrook2024rigorous] was introduced to validate candidate eigenvalues by computing residuals of approximate eigenfunctions. The key insight is that small residuals provide a certificate that a computed eigenvalue lies near the true spectrum, while large residuals indicate potential spectral pollution. This approach achieves rigorous convergence to spectral measures of the Koopman operator and provides a mathematically prominent framework for spectral computation. However, its implementation requires careful estimation of approximation errors associated with finite-dimensional projections, as these errors can be sensitive to the choice of dictionary and to the underlying spectral structure.
@@
In this paper, we explore an alternative approach based on direct construction of the pseudo-resolvent operator from data [giannakis2024consistent, herrmann2021data, kato1995perturbation, susuki2021koopman]. The resolvent encodes complete spectral information through its operator norm, which becomes large precisely when the spectral parameter approaches the spectrum. We show that the Sherman-Morrison-Woodbury identity enables efficient computation of the pseudo-resolvent directly from data matrices, and we extract pseudoeigenfunctions as directions of maximal amplification under this operator. We establish convergence of the approximate spectrum to the true spectrum in the Hausdorff metric with preservation of algebraic multiplicities, and derive computable error bounds for eigenvalue approximations [chaitin1983spectral, kato1995perturbation, 10.1137/0716052]. Numerical experiments on pendulum, Lorenz, and coupled oscillator systems demonstrate the effectiveness of the method in suppressing spectral pollution and resolving closely spaced spectral components.
@@
The remainder of this paper is organized as follows. Section 2 reviews the necessary background on Koopman operators, resolvents, and pseudo-resolvents. Section 3 presents the data-driven construction of the pseudo-resolvent matrix. Section 4 develops the convergence theory, including Hausdorff convergence of spectra and conditions for multiplicity preservation. Section 5 derives error bounds controlled by computable residual quantities. Section 6 presents numerical experiments on pendulum, Lorenz, and coupled oscillator systems. Section 7 concludes with a discussion of the method’s scope and directions for future work.
2 Preliminary
@@ 2.1 Koopman Operator
The Koopman operator is a cornerstone of modern dynamical systems theory, offering a linear framework to study the evolution of nonlinear systems. Consider a dynamical system defined on the state space $\mathcal{M}$ equipped with a measure $\mu$ (which is not necessarily invariant), and let $T:\mathcal{M}\to\mathcal{M}$ be the transformation map. The bounded Koopman operator $\mathcal{K}$ associated with $T$ acts on the space of square-integrable functions $\mathcal{F}\coloneqq L^{2}(\mathcal{M},\mu)$, with the inner product defined by $\langle f,g\rangle\coloneqq\int_{\mathcal{M}}\overline{f}g\,d\mu$, is given by
@@
Under standard assumptions, e.g., $C_{0}$ semigroup [engel1999one, pazy2012semigroups], this Koopman semigroup has a closed and densely defined generator $\mathcal{A}$ such that $\mathcal{K}^{t}=e^{t\mathcal{A}}$ for $t\geq 0$. In our study, the discrete-time operator $\mathcal{K}$ is viewed as the embedded system obtained by sampling the continuous-time flow at unit time $t=1$, i.e.,
@@
This relationship allows us to leverage the rich theoretical framework of discrete approximation for convergence analysis while establishing a direct connection to the continuous-time dynamics. In the pendulum example of Section 6.1, the first and second Figures 1, 2 are based on the discrete operator $\mathcal{K}$ (or its finite-dimensional approximation), whereas the third Figure 3 investigates the continuous-time generator. This dual perspective ensures that our discrete framework not only rigorously captures the spectral properties of $\mathcal{K}$ but also provides a pathway to understand the continuous-time dynamics via this embedding as in Eq. (1).
@@
In practice, the Koopman operator is particularly valuable in data-driven analysis, as it allows one to approximate $\mathcal{K}$ directly from time-series data without requiring an explicit model for $T$. This approach has been successfully applied in diverse fields such as fluid dynamics, neuroscience, and statistical mechanics, where the evolution of observables is often more accessible than tracking individual trajectories.
@@ 2.2 Resolvent and Pseudo-Resolvent Operator
A generalization of the resolvent operator is the pseudo-resolvent, which extends the framework beyond cases where a well-defined generating operator exists. A family of operators $\{\mathcal{R}(z)\}_{z\in\Omega}$, where $\Omega$ is a subset of the complex plane $\mathbb{C}$, is called a pseudo-resolvent if it satisfies the first resolvent identity:
@@ 3 A Data-Driven Computational Approach for Pseudo-Resolvent Estimation
Recall that the EDMD-approximated Koopman operator(semigroup) matrix $\mathbf{K}_{N}$ [Williams_2015] and generator matrix $\mathbf{A}_{N}$ [KLUS2020132416] are:
@@
where $\Psi_{X},\Psi_{Y}\in\mathbb{C}^{M\times N}$ are $N$ basis evaluations at $M$ data points, $\Psi_{X}^{\prime}=(\Psi_{Y}-\Psi_{X})/\Delta t$ and $\Psi_{X}^{\dagger}$ is pseudoinverse of $\Psi_{X}$.
@@
Alternatively, one can also apply Itô formula to compute each elements in the matrix $\Psi_{X}^{\prime}$ if the system is stochastic. See [KLUS2020132416, xu2025datadrivenframeworkkoopmansemigroup] for more detals.
The Sherman-Morrison-Woodbury (SMW) [woodbury1950inverting] identity is the following:
@@
Note that, the construction of pseudo-resolvent matrix $\mathbf{R}_{N}(z)$ is equivalent to the resolvent matrix $\mathbf{R}(z,\mathbf{K}_{N})$ by Eq.(2); specifically,
@@
Thus, by Eq.(5), we can verify matrix identity for difference of inverses in the following:
@@
Analogously, the corresponding pseudo-resolvent matrix $\left(zI-\Psi_{X}^{\dagger}\Psi_{X}^{\prime}\right)^{-1}$ for a generator study in a continuous time setting is given by:
@@
In this paper, we will use notation $\mathbf{R}_{N}(z)$ for both $\mathbf{R}(z,\mathbf{K}_{N})$ and $\mathbf{R}(z,\mathbf{A}_{N})$, depending on the context of discussion.
@@ 4 Convergence Analysis of Spectrum
Denote $\sigma(\mathcal{K})$ the spectrum of bounded Koopman operator $\mathcal{K}$. Let $\mathcal{K}_{N}$ be a sequence of finite-dimensional discretized operators approximating $\mathcal{K}$ due to the discretization by Galerkin approximation [korda2018convergence, Williams_2015]. The resolvent operator $\mathcal{R}(z)$ is defined as:
@@
for $z\in\rho(\mathcal{K})$, where $\rho(\mathcal{K})$ is the resolvent set of $\mathcal{K}$ [kato1995perturbation].
@@
Notice that the convergence of $\lim_{M\to\infty}\mathbf{R}_{N}(z)$ in large data is guaranteed by examining its structure in Eq.(4). This expression can be expanded as a Neumann series where each term involves powers of the Koopman operator matrix. According to EDMD theory [korda2018convergence, Williams_2015], these matrices converge almost surely in the large data limit $M\to\infty$, where $M$ represents the number of data points. Specifically, under appropriate sampling conditions, the empirical averages(i.e., the two Gram matrices formed by Galerkin approximation) in the EDMD procedure converge to their expected values by the Strong Law of Large Numbers(SLLN), ensuring that our pseudo-resolvent operator approximation $\mathcal{R}_{N}(z)$ is well-defined and reliable for spectral analysis. Moreover, since $\mathcal{K}$ is bounded, $\mathcal{K}_{N}\to\mathcal{K}$ in strong operator topology as proved in [korda2018convergence, Theorem 3] However, this does not implies the same convergence type for $\mathcal{R}_{N}(z)\to\mathcal{R}(z)$. It is only true with extra condition, as we will discuss about later.
@@
For any $z\in\rho(\mathcal{K})$ there exists $N(z)$ such that for all $N>N(z)$, we have
@@
for some $M_{z}<\infty$.
Assumption 4.2 is usually called stable convergence [chaitin1983spectral, Chapter 3.4]. It acts as a stability condition that prevents ill-conditioning in the approximation process. Without such a condition, spectral approximation could produce spurious eigenvalues or fail to capture the true spectral properties of the original operator.
@@
Recall second resolvent identity: $\mathcal{R}_{N}(z)-\mathcal{R}(z)=\mathcal{R}_{N}(z)(\mathcal{K}-\mathcal{K}_{N})\mathcal{R}(z)$. By Eq.(7) and the second resolvent identity, we obtain a strong convergence for $\mathcal{R}_{N}(z)$, i.e., for each $z\in U\subset\rho(\mathcal{K})$, there exists $N_{z}\in\mathbb{N}$ such that for $N>N_{z}$,
@@
Assumption 4.5 is usually called strongly stable convergence [chaitin1983spectral, Chapter 5] which is essential for multiplicity preservation in the spectral convergence. This assumption is valid since we will later show that $\text{dim}\,\mathcal{P}_{N}^{\lambda}\mathcal{F}\geq m_{\lambda}$ whenever $N$ is large enough in Corollary 4.9.
@@
First, we show that $\mathcal{R}(z)$ is continuous in $z\in U$ and uniformly in $N$ for $N$ large enough.
Denote $M_{z}=\sup_{N>N_{z}}\|\mathcal{R}_{N}(z)\|<\infty$, which is finite for each $z$. For any $\varepsilon>0$, since $U$ is a compact subset of $\rho(\mathcal{K})$, $U$ can be covered by a finite number $J$ of open balls $\{B(z_{j},\varepsilon)\}_{j=1}^{J}$, where $z_{j}\in U$. Then, for each $z_{j}$, Eq.(7) ensures there exists $N_{z_{j}}$ such that for $N>N_{z_{j}}$, $\|\mathcal{R}_{N}(z_{j})\|\leq M_{z_{j}}<\infty$. Define $N_{0}=\max_{j=1,\ldots,J}N_{z_{j}}$ and $M_{0}=\max_{j=1,\ldots,J}M_{z_{j}}$. Thus, for $N>N_{0}$, $\|\mathcal{R}_{N}(z_{j})\|\leq M_{0}$ at these finitely many points.
For any $z\in K$, there exists $z_{j}$ such that $|z-z_{j}|<\varepsilon$. Using the first resolvent identity:
$\mathcal{R}_{N}(z)-\mathcal{R}_{N}(z_{j})=\mathcal{R}_{N}(z)(z-z_{j})\mathcal{R}_{N}(z_{j}),$
we obtain:
$\|\mathcal{R}_{N}(z)-\mathcal{R}_{N}(z_{j})\|\leq\|\mathcal{R}_{N}(z)\|\|\mathcal{R}_{N}(z_{j})\||z-z_{j}|<\|\mathcal{R}_{N}(z)\|M_{0}\varepsilon.$
So, we have shown that $\mathcal{R}(z)$ is continuous in $z\in U$, and thus $\|\mathcal{R}(z)\|$ is bounded on $U$. Next,
$\|\mathcal{R}_{N}(z)\|\leq\|\mathcal{R}_{N}(z)-\mathcal{R}_{N}(z_{j})\|+\|\mathcal{R}_{N}(z_{j})\|\leq\|\mathcal{R}_{N}(z)\|M_{0}\varepsilon+M_{0}.$
After rearranging, we have:
$\|\mathcal{R}_{N}(z)\|(1-M_{0}\varepsilon)\leq M_{0}.$
Choose $\varepsilon>0$ such that $M_{0}\varepsilon<1$ (e.g., $\varepsilon<1/(M_{0}+1)$), ensuring $1-M_{0}\varepsilon>0$. Then
$\|\mathcal{R}_{N}(z)\|\leq\frac{M_{0}}{1-M_{0}\varepsilon}.$
Let $M_{U}\coloneqq\frac{M_{0}}{1-M_{0}\varepsilon}$, which depends only on $U$, we finish the proof. ∎
Lemma 4.7 ([chaitin1983spectral, Proposition 3.10]).
Let $\mathcal{P}$ and $\mathcal{P}_{N}$ (for $n\in\mathbb{N}$) be projections such that $\mathcal{P}_{N}$ converges to $P$ pointwise, and let $\dim\mathcal{P}\mathcal{F}<\infty$. Then, $\|(\mathcal{P}-\mathcal{P}_{N})\mathcal{P}\|\to 0$ as $N$ becomes sufficiently large, and for $N$ large enough, $\dim\mathcal{P}_{N}\mathcal{F}\geq\dim\mathcal{P}\mathcal{F}$.
@@
In many numerical methods, particularly those that rely on projecting infinite-dimensional problems onto finite-dimensional subspaces (e.g. Galerkin approximation), Lemma 4.7 plays a fundamental role in ensuring stability and consistency of approximations. It guarantees that once $\mathcal{P}_{N}$ converges strongly to a limiting projector $\mathcal{P}$, the dimension of the projected subspace $\mathcal{P}_{N}\mathcal{F}$ cannot suddenly drop below the dimension of $\mathcal{P}_{N}\mathcal{F}$. This means that the subspace capturing the “essential” part of the solution does not degenerate.
@@
Suppose Eq.(7) in Assumption 4.2 holds for all $z\in\Gamma$. Then, $\text{dim}\,\mathcal{P}_{N}\mathcal{F}\geq\text{dim}\,\mathcal{P}\mathcal{F}$, for $N$ large enough.

### v2

@@ Abstract
We present a data-driven method for spectral analysis of the Koopman operator based on direct construction of the pseudo-resolvent from time-series data. Finite-dimensional approximations of the Koopman operator, such as those obtained from Extended Dynamic Mode Decomposition, are known to suffer from spectral pollution. To address this issue, we construct the pseudo-resolvent operator using the Sherman-Morrison-Woodbury identity whose norm serves as a spectral indicator, and pseudoeigenfunctions are extracted as directions of maximal amplification. We establish convergence of the approximate spectrum to the true spectrum in the Hausdorff metric for isolated eigenvalues, with preservation of algebraic multiplicities, and derive error bounds for eigenvalue approximation. Numerical experiments on pendulum, Lorenz, and coupled oscillator systems demonstrate that the method effectively suppresses spectral pollution and resolves closely spaced spectral components.
@@ 1 Introduction
The study of nonlinear dynamical systems is central to many scientific and engineering problems [1, 2, 15, 27, 30, 31]. Traditionally, their analysis relies on first-principles models together with tools such as phase-space geometry [27], perturbation methods [14] for weak nonlinearities, and direct numerical simulation. These approaches have delivered deep insight, but they can be difficult to use when an accurate governing model is unavailable, when model parameters are uncertain, or when the goal is to infer global dynamical features from observations rather than describe local behavior near equilibria [10, 11]. Meanwhile, modern experiments and simulations increasingly produce large amounts of time-series measurements, which has attracted growing interest in data-driven methods that aim to characterize dynamical behavior directly from data.
@@
A natural perspective for data-driven analysis is to focus on the evolution of observable quantities rather than the underlying state variables. In many applications, one does not have direct access to the full state of a system but instead observations of the state, e.g., temperatures, velocities, concentrations, or other quantities of interest. The Koopman operator provides a mathematical framework for this viewpoint; more specifically, it describes how observables evolve in time by composing them with the flow map of the dynamical system [19, 18, 23]. Although the underlying dynamics may be nonlinear, the Koopman operator acts linearly on the space of observables, enabling spectral analysis to reveal dynamical features such as oscillation frequencies, growth and decay rates, and coherent structures. Dynamic Mode Decomposition (DMD) [33] and its extension, Extended Dynamic Mode Decomposition (EDMD) [34], provide practical algorithms for approximating the Koopman operator and its spectral properties directly from time-series data, and theoretical foundations have established convergence of these approximations under appropriate conditions [5, 7, 13, 17, 20, 36].
@@
Despite these convergence guarantees, a fundamental difficulty arises when extracting spectral information from finite-dimensional approximations: the phenomenon of spectral pollution [5]. The approximating matrix is finite-dimensional and therefore possesses a discrete spectrum, whereas the true Koopman operator includes continuous spectral components (see, e.g., [22]). Consequently, computed eigenvalues may include spurious points that do not correspond to any part of the true spectrum. In other words, these are drawbacks of the finite-dimensional truncation rather than intrinsic dynamical features. Conversely, true spectral points may be missed entirely if they are not well-represented by the chosen dictionary. This problem is not merely a matter of insufficient data or dictionary size; it is a structural consequence of approximating an infinite-dimensional operator by finite-dimensional matrices.
@@
To address spectral pollution, Residual Dynamic Mode Decomposition [5] was introduced to validate candidate eigenvalues by computing residuals of approximate eigenfunctions. The key insight is that small residuals provide a certificate that a computed eigenvalue lies near the true spectrum, while large residuals indicate potential spectral pollution. This approach achieves rigorous convergence to spectral measures of the Koopman operator and provides a rigorous framework for spectral computation. However, its implementation requires careful estimation of approximation errors associated with finite-dimensional projections, as these errors can be sensitive to the choice of dictionary and to the underlying spectral structure.
@@
In this paper, we explore an alternative approach based on direct construction of the pseudo-resolvent operator from data [9, 12, 14, 32]. The resolvent encodes complete spectral information through its operator norm, which becomes large precisely when the spectral parameter approaches the spectrum. We show that the Sherman-Morrison-Woodbury identity enables efficient computation of the pseudo-resolvent directly from data matrices, and we extract pseudoeigenfunctions as directions of maximal amplification under this operator. We establish convergence of the approximate spectrum to the true spectrum in the Hausdorff metric with preservation of algebraic multiplicities, and derive error bounds for eigenvalue approximations [4, 14, 25]. Numerical experiments on pendulum, Lorenz, and coupled oscillator systems demonstrate the effectiveness of the method in suppressing spectral pollution and resolving closely spaced spectral components.
@@
The remainder of this paper is organized as follows. Section 2 reviews the necessary background on Koopman operators, resolvents, and pseudo-resolvents. Section 3 presents the data-driven construction of the pseudo-resolvent matrix. Section 4 develops the convergence theory, including Hausdorff convergence of spectra and conditions for multiplicity preservation. Section 5 derives error bounds controlled by residual quantities. Section 6 presents numerical experiments on pendulum, Lorenz, and coupled oscillator systems. Section 7 concludes with a discussion of the method’s scope and directions for future work.
2 Preliminaries
@@ 2.1 Koopman Operator
The Koopman operator is a cornerstone of modern dynamical systems theory, offering a linear framework to study the evolution of nonlinear systems. Consider a dynamical system defined on the state space $\mathcal{M}$ equipped with a measure $\mu$ (which is not necessarily invariant), and let $T:\mathcal{M}\to\mathcal{M}$ be the transformation map. The bounded Koopman operator $\mathcal{K}$ associated with $T$ acts on the space of square-integrable functions $\mathcal{F}\coloneqq L^{2}(\mathcal{M},\mu)$. The inner product is defined by $\langle f,g\rangle\coloneqq\int_{\mathcal{M}}\overline{f}g\,d\mu$, and the operator is given by
@@
Under standard assumptions, e.g., $C_{0}$ semigroup [8, 26], this Koopman semigroup has a closed and densely defined generator $\mathcal{A}$ such that $\mathcal{K}^{t}=e^{t\mathcal{A}}$ for $t\geq 0$. In our study, the discrete-time operator $\mathcal{K}$ is viewed as the embedded system obtained by sampling the continuous-time flow at unit time $t=1$, i.e.,
@@
This relationship allows us to leverage the rich theoretical framework of discrete approximation for convergence analysis while establishing a direct connection to the continuous-time dynamics. In the pendulum example of Section 6.1, Figures 1 and 2 are based on the discrete operator $\mathcal{K}$ (or its finite-dimensional approximation), whereas Figure 3 investigates the continuous-time generator. This dual perspective ensures that our discrete framework not only rigorously captures the spectral properties of $\mathcal{K}$ but also provides a pathway to understand the continuous-time dynamics via this embedding as in Eq. (1).
@@
In practice, the Koopman operator is particularly valuable in data-driven analysis [3, 5, 13, 17, 21, 23, 24, 34, 36, 37], as it allows one to approximate $\mathcal{K}$ directly from time-series data without requiring an explicit model for $T$. This approach has been successfully applied in diverse fields such as fluid dynamics, neuroscience, and statistical mechanics, where the evolution of observables is often more accessible than tracking individual trajectories.
@@ 2.2 Resolvent and Pseudo-Resolvent Operator
A generalization of the resolvent operator is the pseudo-resolvent, which extends the framework beyond cases where a well-defined generating operator exists. A family of operators $\{\mathcal{R}(z)\}_{z\in\Omega}$, where $\Omega$ is a subset of the complex plane $\mathbb{C}$, is called a pseudo-resolvent if it satisfies the first resolvent identity [8, 14]:
@@ 3 A Data-Driven Computational Approach for Pseudo-Resolvent Estimation
Recall that the EDMD-approximated Koopman operator (semigroup) matrix $\mathbf{K}_{N}$ [34] and generator matrix $\mathbf{A}_{N}$ [17] are:
@@
where $\Psi_{X},\Psi_{Y}\in\mathbb{C}^{M\times N}$ are $N$ basis evaluations at $M$ data points, $\Psi_{X}^{\prime}=(\Psi_{Y}-\Psi_{X})/\Delta t$ and $\Psi_{X}^{\dagger}$ is the pseudoinverse of $\Psi_{X}$.
@@
Alternatively, one can also apply the Itô formula to compute each element in the matrix $\Psi_{X}^{\prime}$ if the system is stochastic. See [17, 36] for more details.
The Sherman-Morrison-Woodbury (SMW) [35] identity is the following:
@@
Note that the construction of pseudo-resolvent matrix $\mathbf{R}_{N}(z)$ is equivalent to the resolvent matrix $\mathbf{R}(z,\mathbf{K}_{N})$ by Eq. (2); specifically,
@@
Thus, by Eq. (5), we can verify the matrix identity for difference of inverses in the following:
@@
Analogously, the corresponding pseudo-resolvent matrix $\left(zI-\Psi_{X}^{\dagger}\Psi_{X}^{\prime}\right)^{-1}$ for a generator study in a continuous-time setting is given by:
@@
In this paper, we will use the notation $\mathbf{R}_{N}(z)$ for both $\mathbf{R}(z,\mathbf{K}_{N})$ and $\mathbf{R}(z,\mathbf{A}_{N})$, depending on the context of the discussion.
@@
Remark 3.3.
Residual DMD [5] and the resolvent-based approach developed here are closely related in that both aim to mitigate spectral pollution by accessing pseudospectral information of the Koopman operator. Conceptually, these are two complementary viewpoints of the same underlying quantity: the $\varepsilon$-pseudospectrum can be characterized either through a residual condition (existence of a unit-norm vector with small defect) or through resolvent amplification. ResDMD adopts the residual viewpoint and provides a posteriori validation for candidate spectral points by computing minimal residuals of approximate eigenfunctions (and related adjoint information). In contrast, our method constructs a data-driven approximation of the (pseudo-)resolvent directly and extracts pseudospectral features through its norm and dominant amplification directions. From a practical standpoint, residual-based certification can be advantageous when the goal is to validate a small number of candidate eigenpairs, whereas resolvent-based probing naturally produces global pseudospectral contour plots and associated pseudomodes, at the cost of evaluating the resolvent over a set of spectral parameters.
@@ 4 Convergence Analysis of Spectrum
Denote by $\sigma(\mathcal{K})$ the spectrum of the bounded Koopman operator $\mathcal{K}$. Let $\mathcal{K}_{N}$ be a sequence of finite-dimensional discretized operators approximating $\mathcal{K}$ resulting from the discretization by Galerkin approximation [20, 34]. The resolvent operator $\mathcal{R}(z)$ is defined as:
@@
for $z\in\rho(\mathcal{K})$, where $\rho(\mathcal{K})$ is the resolvent set of $\mathcal{K}$ [14].
@@
Notice that the convergence of $\lim_{M\to\infty}\mathbf{R}_{N}(z)$ in large data is guaranteed by examining its structure in Eq. (4). This expression can be expanded as a Neumann series where each term involves powers of the Koopman operator matrix. According to EDMD theory [20, 34], these matrices converge almost surely in the large data limit $M\to\infty$, where $M$ represents the number of data points. Specifically, under appropriate sampling conditions, the empirical averages (i.e., the two Gram matrices formed by Galerkin approximation) in the EDMD procedure converge to their expected values by the Strong Law of Large Numbers (SLLN), ensuring that our pseudo-resolvent operator approximation $\mathcal{R}_{N}(z)$ is well-defined and reliable for spectral analysis. Moreover, since $\mathcal{K}$ is bounded, $\mathcal{K}_{N}\to\mathcal{K}$ in strong operator topology as proved in [20, Theorem 3]. However, this does not imply the same convergence type for $\mathcal{R}_{N}(z)\to\mathcal{R}(z)$. It is only true under an extra condition, as we will discuss later.
@@
For any $z\in\rho(\mathcal{K})$ there exists $N(z)$ such that for all $N>N(z)$, we have $z\in\rho(\mathcal{K}_{N})$ and
@@
for some $M_{z}<\infty$ independent of $N$.
Assumption 4.2 is usually called stable convergence [4, Chapter 3.4]. It acts as a stability condition that prevents ill-conditioning in the approximation process. Without such a condition, spectral approximation could produce spurious eigenvalues or fail to capture the true spectral properties of the original operator.
@@
Recall second resolvent identity [8, 14]: $\mathcal{R}(z)-\mathcal{R}_{N}(z)=\mathcal{R}_{N}(z)(\mathcal{K}-\mathcal{K}_{N})\mathcal{R}(z)$. By Eq. (7) and the second resolvent identity, we obtain a strong convergence for $\mathcal{R}_{N}(z)$, i.e., for each $z\in U\subset\rho(\mathcal{K})$, there exists $N_{z}\in\mathbb{N}$ such that for $N>N_{z}$,
@@
Assumption 4.5 is usually called strongly stable convergence [4, Chapter 5] which is essential for multiplicity preservation in the spectral convergence. This assumption is valid since we will later show that $\text{dim}\,\mathcal{P}_{N}^{\lambda}\mathcal{F}\geq m_{\lambda}$ whenever $N$ is large enough in Corollary 4.9.
@@
Fix any $z\in U$. By Assumption 4.2, there exist $N(z)$ and $M_{z}<\infty$ such that for all $N>N(z)$, $z\in\rho(\mathcal{K}_{N})$ and $\|\mathcal{R}_{N}(z)\|\leq M_{z}$. Define $r_{z}:=\frac{1}{2M_{z}}$.
Now take any $w\in\mathbb{C}$ with $|w-z|<r_{z}$ and any $N>N(z)$. Using the first resolvent identity,
$\mathcal{R}_{N}(w)-\mathcal{R}_{N}(z)=(z-w)\mathcal{R}_{N}(w)\mathcal{R}_{N}(z),$
we obtain
$\mathcal{R}_{N}(w)\bigl[I+(w-z)\mathcal{R}_{N}(z)\bigr]=\mathcal{R}_{N}(z).$
Since $\|(w-z)\mathcal{R}_{N}(z)\|\leq|w-z|\,\|\mathcal{R}_{N}(z)\|<\frac{1}{2}$, the operator $I+(w-z)\mathcal{R}_{N}(z)$ is invertible and
$\|\mathcal{R}_{N}(w)\|\leq\frac{\|\mathcal{R}_{N}(z)\|}{1-|w-z|\,\|\mathcal{R}_{N}(z)\|}\leq 2M_{z}.$
Hence,
$\sup_{w\in B(z,r_{z}),\,N>N(z)}\|\mathcal{R}_{N}(w)\|\leq 2M_{z}.$
The collection $\{B(z,r_{z})\}_{z\in U}$ is an open cover of the compact set $U$, so we can extract a finite subcover $\{B(z_{j},r_{z_{j}})\}_{j=1}^{J}$. Let
$N_{U}:=\max_{1\leq j\leq J}N(z_{j}),\qquad M_{U}:=2\max_{1\leq j\leq J}M_{z_{j}}.$
Then for any $N>N_{U}$ and any $w\in U$, there exists $j$ with $w\in B(z_{j},r_{z_{j}})$ and therefore $\|\mathcal{R}_{N}(w)\|\leq M_{U}$. This proves the claim. ∎
Lemma 4.7 ([4, Proposition 3.10]).
Let $\mathcal{P}$ and $\mathcal{P}_{N}$ (for $n\in\mathbb{N}$) be projections such that $\mathcal{P}_{N}$ converges to $\mathcal{P}$ pointwise, and let $\dim\mathcal{P}\mathcal{F}<\infty$. Then, $\|(\mathcal{P}-\mathcal{P}_{N})\mathcal{P}\|\to 0$ as $N$ becomes sufficiently large, and for $N$ large enough, $\dim\mathcal{P}_{N}\mathcal{F}\geq\dim\mathcal{P}\mathcal{F}$.
@@
In many numerical methods, particularly those that rely on projecting infinite dimensional problems onto finite-dimensional subspaces (e.g. Galerkin approximation), Lemma 4.7 plays a fundamental role in ensuring stability and consistency of approximations. It guarantees that once $\mathcal{P}_{N}$ converges strongly to a limiting projector $\mathcal{P}$, the dimension of the projected subspace $\mathcal{P}_{N}\mathcal{F}$ cannot suddenly drop below $\dim(\mathcal{P}\mathcal{F})$. This means that the subspace capturing the “essential” part of the solution does not degenerate.
@@
Let $\lambda$ be an isolated eigenvalue of $\mathcal{K}$ with algebraic multiplicity $m_{\lambda}$ and let $\Gamma_{\lambda}$ be the isolating contour in Definition 4.4. Suppose Eq. (7) in Assumption 4.2 holds for all $z\in\Gamma_{\lambda}$. Then, for $N$ sufficiently large,
$\dim\bigl(\mathcal{P}_{N}^{\lambda}\mathcal{F}\bigr)\geq\dim\bigl(\mathcal{P}^{\lambda}\mathcal{F}\bigr)=m_{\lambda}.$